The following table shows the mean number of police brutality complaints made to Louisiana State Police Internal Affairs per month before and after the institution of body cameras on police officers. Test the significance of the difference in the mean number of police brutality complaints made to Louisiana State Police Internal Affairs before and after the institution of body cameras on police officers (α=0.001). Show all details of the test.

Before Body Cameras 24, 15, 34, 9, 28, 44

After Body Cameras 24 20 18 29 8 24 38
Answer #1

X,: The number of police brutality complaints made to Lauisiana state police Internal Affairs per month before the institutioJ di - 17 J = 2.8933 d = 2.8333 sd = Eld- 2)2 = 54.8333 - 154.8333 sd = sd - 10.97 3.31.16 Ho: eldo vs Hilld to mean differenP-value = 0.09021 Decision Here we Refect Ho if p-valve 2 = 0.001 pevale =0.ogo2 t = 0.001 failed to reject Ho Conclusion we